### v1

# Eco-evolutionary constraints for the endemicity of rapidly evolving viruses

Abstract
Antigenic escape constitutes the main mechanism allowing rapidly evolving viruses to achieve endemicity. Beyond granting immune escape, empirical evidence also suggests that mutations of viruses might increase their inter-host transmissibility. While both mechanisms are well-studied individually, their combined effects on viral endemicity remain to be explored. Here we propose a minimal eco-evolutionary framework to simulate epidemic outbreaks generated by pathogens evolving both their transmissibility and immune escape. Our findings uncover a very rich phenomenology arising from the complex interplay between both evolutionary pathways and the underlying contagion dynamics. We first show that contagions at the population level constrain the effective evolution of the virus, accelerating the increase in transmissibility in the first epidemic wave while favoring antigenic variation in the transition to the endemic phase. Our results also reveal that accounting for both evolutionary pathways changes the features of the viruses more prone to become endemic. While chances for endemicity increase with infectiousness of the wild-type variant for viruses not evolving their transmissibility, a non-monotonic behavior is observed when the latter mechanism is included, favoring less transmissible viruses and impairing those ones with intermediate infectiousness.
Introduction
Modeling the propagation of rapidly evolving viruses poses a major theoretical challenge given the compatibility between the evolutionary and epidemiological time scales [1]. Indeed, the evolution of viruses throughout epidemic outbreaks increases the richness of epidemic trajectories and renders complex and fast-changing variants landscapes [2, 3] such as the one recently observed for the SARS-CoV-2 virus [4]. Hence, modeling mathematically these intricate dynamics requires moving from the classical compartmental models [5] to eco-evolutionary frameworks [6, 7], explicitly accounting for the rate at which mutations occur, their effects on virus’ fitness and the ecological dynamics generated by the competition of the multiple co-circulating strains in the same population.
Yet eliciting an immune response in their hosts, some rapidly evolving viruses persist in real populations due to the antigenic variation arising from mutations in antibody-binding regions [8, 9, 10, 11]. This evolutionary mechanism facilitates reinfection events by antigenically distant variants that evade pre-existing immunity within the population. The relation between reinfection events and antigenic distances has been incorporated into theoretical models, assuming that variants can be embedded as nodes in genotype networks [12, 13, 14] or in discrete [15, 16, 17, 18, 17] or continuous [19, 20] low-dimensional representations of the antigenic space. Rouzine et al. [19] show that the interplay between immune presure and virus evolution yields travelling fitness waves in 1D antigenic spaces, sustaining enough viral diversity for the virus to persist in the population. Morever, antigenic spaces with higher dimensionality can also harbor the speciation of viruses, splitting at some point into different sublineages which evolve independently [20, 18].
Pathogens’ evolution also shapes other non-antigenic traits such as their inter-host transmissibility [21, 22] or their virulence [23, 24], leading to major changes in their associated epidemic trajectories [25, 26, 27] and their impact on public health [28, 29]. For instance, recent genomic surveillance data reveals that SARS-CoV-2 virus has undergone both evolution in its transmissibility and antigenic escape [30, 31]. Motivated by this empirical evidence, several theoretical works have proven that the underlying eco-evolutionary dynamics for these viruses intertwine both evolutionary pathways. Namely, non-antigenic traits can evolve towards maximizing either the basic reproduction number ${\mathcal{R}_{0}}$ or the speed of the antigenic fitness wave at equilibrium as a function of the immune pressure existing in the population [32, 33].
As the SARS-CoV-2 virus transitions from an epidemic to an endemic phase [34], there is increasing interest in determining the eco-evolutionary constraints that enable this transition in real populations [35, 36]. To the best of our knowledge, understanding how evolution in both antigenic and non-antigenic traits determines the endemicity of viruses remains an open question. To fill this gap, here we propose a minimal eco-evolutionary framework, extending the classical Susceptible-Infected-Recovered (SIR) model to integrate the evolution of pathogens in both their transmissibility and their immune escape. In absence of the former evolutionary pathway, we show that the probability of a virus becoming endemic depends on its antigenic escape and always increases with the infectiousness of the wild-type variant. In contrast, our stochastic simulations reveal that evolution in non-antigenic traits significantly alter this picture, resulting in a non-monotonic behavior where viruses with low infectiousness display greater endemicity than others with intermediate infectiousness. We derive a heuristic equation capturing this behavior as a result of the trade-off between the minimum infectiousness needed to sustain an endemic scenario and the characteristic evolutionary to reach such infectiousness. Finally, we discuss the practical implications of our findings.
Results
.1 Model for the coevolution of transmissibility and antigenic escape
Fig. 1 sketches the main features of the minimal eco-evolutionary framework here introduced. We assume a constant and well-mixed population of $N$ individuals making $k$ contacts per time step. As stated above, from an epidemiological point of view, our framework is a modified version of the SIR model to harbor reinfection events for those variants escaping the immune response mounted in the population. In this sense, each virus $j$ circulating across the population is characterized by two variables: its transmissibility ${\lambda_{j}}$ and its position in the antigenic space $x_{j}$. For the sake of simplicity, we consider a 1D antigenic space, aligning with previous theoretical frameworks [33, 19].
Fig. 1A illustrates all the processes changing the epidemiological state of individuals in our model. Namely, each infected individual $j$ recover at a rate $\mu$, preserving the antigenic position $x_{j}$ of the variant for which they developed their immune response. Regarding contagions, we assume that any susceptible individual in contact with an infected individual $j$ contracts the pathogen at a rate $\lambda_{j}$. Conversely, the rate of reinfection events $\lambda^{\prime}_{ij}$ is different for each recovered individual $i$, as it should account for the antigenic distance between its associated variant and that of the infected agent. Following [15, 19], for each pair $(i,j)$ formed by a recovered individual $i$ interacting with an infected agent $j$, we assume that the contagion rate is given by:
$\lambda^{\prime}_{ij}=\begin{cases}\lambda_{j}\left(1-e^{-(x_{j}-x_{i})}\right)&\text{if }x_{j}\geq x_{i}\\ 0&\text{otherwise}\end{cases}$ (1)
The latter choice imposes that reinfection events only occurs when $x_{j}>x_{i}$. Such assumption gives rise a travelling wave of variants moving with a positive speed in the 1D antigenic space [19]. Hence, our model assumes that natural immunity against past variants does not wane over time. For all contagions, the agent contracting the pathogen inherits the variant of the infected individual.
Figure 1: Schematic of the eco-evolutionary model here introduced. A) Modified SIR model to account for the reinfection events driven by the immune escape of the virus. Each infected agent $j$ recovers at a rate $\mu$, keeping their associated antigenic position $x?j$. Primary infections caused by an infected agent $j$ occur at a rate $\lambda_{j}$, whereas the contagion rate for a recovered individual $i$ in contact with that infected individual is $\lambda^{\prime}_{ij}=\Theta(x_{j}-x_{i})\lambda_{j}\left(1-e^{-(x_{j}-x_{i})}\right)$, where $\Theta(x)$ represents the Heavyside function. Note that the latter increases with the antigenic distance between variants and that reinfection is only possible when $x_{j}>x_{i}$, thus assuming that natural immunity against a variant does not wane over time. B) Evolution of the virus transmissibility and antigenic position in infected individuals, assuming that changes in trait $m$ ($m\in\left\{x,\lambda\right\}$) follows $\Delta_{m}\sim\mathcal{N}(0,D^{2}_{m})$, with $D^{2}_{m}$ determining its speed of evolution.
Fig. 1B shows the evolutionary processes changing the epidemiological parameters of the circulating viral strains. We assume that mutations just occur in infected individuals, shaping both the transmissibility of the virus and its antigenic position by a factor $\Delta\lambda$ and $\Delta x$ respectively. Inspired by previous works [25, 19, 32], we assume that changes in transmissibility (antigenic position) are drawn from a zero-mean Gaussian distribution whose variance $D^{2}_{\lambda}$ ($D^{2}_{x}$) determines the speed of evolution. Regarding recovered individuals, we assume their antigenic variant to remain constant over time until they get reinfected by another circulating strain. A complete description of the stochastic simulations performed to obtain the epidemic trajectories following this model is provided in the Materials and Methods section.
.2 Endemicity of viruses without transmissibility evolution
To understand the transition from the epidemic to the endemic phase, let us first neglect the evolution of the virus in the transmissibility space by setting $D_{\lambda}=0$ and simulate different epidemic outbreaks varying the speed of evolution in the antigenic space $D_{x}$. Unless otherwise stated, throughout the manuscript we fix the recovery rate of the disease to $\mu=1.0/7$ days-1 and consider epidemic outbreaks spreading across populations of $N=10^{4}$ individuals, each one making $k=10$ contacts per day.
Fig. 2A depicts the three different epidemic regimes which can be found in the model for a pathogen whose wild-type variant has a basic reproduction number $\mathcal{R}^{wt}_{0}=\lambda k/\mu=3$. Recall that the basic reproduction number quantifies the number of contagions triggered by one infected individual in a healthy population. Therefore, when $\mathcal{R}^{wt}_{0}>1$, the pathogen always trigger a first epidemic wave in the population. Nonetheless, the long-term behavior of the epidemic trajectories does depend on the immune escape mechanism, parametrized with $D_{x}$. In absence of immune escape, i.e. $D_{x}=0$, the model reduces to the standard SIR model characterized by a single epidemic wave followed by the extinction of the infected population. Conversely, for $D_{x}=0.03$, a damped oscillatory behavior towards an endemic equilibrium appears for all epidemic trajectories. In between these two extremes ($D_{x}=0.015$), richer epidemic dynamics are generated by the model, as the first epidemic wave is typically followed by long periods of very low epidemic incidence. Such periods act as epidemic bottlenecks, leading to the extinction of some outbreaks due to stochastic fluctuations. Others, however, are able to persist in the population as a result of the enlargement of the distribution of variants in the antigenic space, as shown in Fig. 2B. Comparing the endemic realizations, higher $D_{x}$ values lead to endemic scenarios with a more acute prevalence of the diseases, as reported in [19].
Figure 2: Epidemic trajectories for viruses not evolving their transmissibility. A) Time dependence of the fraction of population in the infected state $\rho_{I}$ as a function of the speed of evolution in the antigenic space $D_{x}$ (color code). B): Time evolution of the variance of the distribution of variants in the antigenic space $\sigma_{x}^{2}$ as a function of $D_{x}$ (color code). In both panels, thick solid (dashed) lines shows the values for those viruses becoming (not becoming) endemic in the population obtained by averaging 200 epidemic outbreaks whereas thin lines represent a sample of 20 individual trajectories in both cases. In these panels, the basic reproduction number of the wild-type variant of the pathogen is set to $\mathcal{R}^{wt}_{0}$=3. C): Fraction of epidemic outbreaks surviving in the population after $t=1000$ days as a function of ${\mathcal{R}^{wt}_{0}}$ and $D_{x}$ (color code). The results shown in this panel have been obtained by simulating $1000$ epidemic outbreaks for each pair of ($\mathcal{R}^{wt}_{0}$,$D_{x}$) values. In all panels, we assume $N=10^{4}$ individuals, $I_{0}=10$ initially infected agents, and we set the recovery rate to $\mu=1/7$ days-1 and the number of daily contacts to $k=10$ interactions.
To further deepen into the features shaping virus’ endemicity, Fig. 2C shows the fraction of epidemic outbreaks surviving in the population after $t=1000$ days as a function of both ${\mathcal{R}_{0}}^{wt}$ and $D_{x}$. There, the three aforementioned epidemic regimes become more evident. Moreover, virus endemicity increases monotonically with the infectiousness of the wild-type variant regardless of the speed of evolution in the antigenic space $D_{x}$. Likewise, for a given $\mathcal{R}_{0}^{wt}$ value, accelerating the evolution in the antigenic space, i.e. increasing $D_{x}$, also enhances virus endemicity, making immune escape more likely.
.3 Endemicity of viruses with transmissibility evolution
So far, we have neglected the evolution of virus transmissibility, retrieving the three dynamical regimes obtained in the literature for the SIR model with immune escape and the expected monotonic behavior of endemicity with virus infectiousness. Hereafter, we include such evolutionary pathway, questioning how the joint evolution of both antigenic and non-antigenic traits alters this picture. We first fix the evolution speed of the different traits by setting $D_{x}=0.015$ and $D_{\lambda}=3\cdot 10^{-4}$ and study the epidemic trajectories for three different pathogens differing in the basic reproduction number of their wild type variants ${\cal R}_{0}^{wt}$.
Fig. 3A shows that evolution in transmissibility allows viruses with low initial infectiousness, e.g. ${\cal R}_{0}^{wt}=1.25$ or ${\cal R}_{0}^{wt}=2$, to become endemic in the population. Note that this result comes in stark contrast with those reported in Fig. 2C, as no endemic realization is observed for this combination of parameters when just accounting for the evolution in antigenic traits. Regarding the shape of the temporal trajectories, the early stages of the outbreak qualitatively resemble the ones shown in Fig. 2A with the presence of a first prominent epidemic wave followed by an epidemic bottleneck. Conversely, the late stages of the outbreak are not characterized by a fixed prevalence of the disease as in the former case. Interestingly, regardless of the infectiousness of the wild-type variant, the epidemic prevalence seems to grow at a (roughly) constant pace on average once the virus has entered the endemic phase.
To shed more light into the impact of the coexistence of both evolutionary pathways on the epidemic trajectories, let us now analyze how each virus trait evolves over time. Fig. 3B shows the time evolution of the associated basic reproduction number ${\cal R}_{0}$ for each pathogen here considered. In all cases, we can observe how transmissibility evolves quicker at early than at late stages of the outbreak. This phenomenon becomes more pronounced when considering viruses with low initial infectiousness and high $D_{\lambda}$ values (see SI Appendix, Fig. S1). Consequently, our results reveal how the underlying eco-evolutionary dynamics yields an accelerated evolution in transmissibility for innocuous pathogens ($\mathcal{R}_{0}^{wt}=1.25$) compared to that experienced by more infectious pathogens, e.g. those ones with ${\cal R}_{0}^{wt}=3$.
[…]
The evolution of antigenic traits can be characterized by studying the distribution of variants in the population across the antigenic space. In contrast with the evolution in transmissibility, we observe in Fig 3C how evolution in the antigenic space mainly occurs in the transition to the endemic regime after the epidemic bottleneck, reflected by the sudden enlargement in the variants distribution observed for the three pathogens. Comparing the different curves, we notice the lower the infectiousness, the higher the peak of antigenic diversity reached by the pathogen. The former result is intuitive, as those pathogens with less infectiousness should escape the immunity existing in the population more efficiently to survive in the long-term. Note that coupling both evolutionary pathways also has strong implications for the long-term evolution of the antigenic diversity. Specifically, antigenic diversity is not stabilized after the initial increase, as in Fig. 2B, but is partially lost as the virus gets more infectious to eventually reach a stable but lower value.
To shed more light into the emergent relationship between both evolutionary pathways, we compute in the Materials and Methods section an approximation for the case reproduction number ${\cal R}^{app}_{case}(t)$ representing the expected number of contagions made by one individual infected at a given day $t$ [37]. The time evolution of this quantity (Fig. 3D) confirms how antigenic and non-antigenic traits compensate one another to produce the same steady growth in the epidemic curves at late stages of the outbreak. This confirms that, despite the absence of any biological trade-off, the combination of the immune pressure in the antigenic space and the evolution in transmissibility yields universal trajectories for the evolution of pathogens in our model.
The analysis of the epidemic trajectories has revealed interesting phenomena arising from integrating evolution of antigenic and non-antigenic traits in a single formalism. In what follows, we focus on our primary research question and study the eco-evolutionary constraints for the endemicity of evolving viruses. Interestingly, Fig. 3E reveals that evolution in transmissibility drastically changes the picture observed in its absence (Fig. 2C). As stated before, evolution in transmissibility promotes virus endemicity, as the fraction of surviving realizations is always higher than in absence of such evolutionary pathway. More strikingly, in situations with strong immune pressure, i.e. low $D_{x}$ values, evolution in transmissibility gives rise to a non-mononotic behavior of the virus endemicity with the infectiousness of the wild-type variant. For instance, the curve corresponding to $D_{x}=0.015$ shows that viruses with an initial reproduction number of $\mathcal{R}_{0}^{wt}\simeq 1$ are more prone to persist in the population than others with $\mathcal{R}_{0}^{wt}\simeq 2$. Note that this unexpected behavior dissappears when immune pressure loses relevance, i.e. for high $D_{x}$ values or high $D_{\lambda}$ values (see SI Appendix, Fig. S1D), retrieving the monotonic behavior typically reported in the literature.
For the sake of completeness, we also consider a scenario where the virus evolution is bounded by a transmission-recovery trade-off accelerating the recovery rate of hosts for more transmissible pathogens (see SI Appendix, Fig. S2). Biologically, the latter assumption is justified as an enhanced immune response for pathogens replicating more efficiently within the host [38]. The results here reported remain robust as long as the relative increase in the recovery rate is negligible. Nonetheless, when the trade-off gains relevance, the non-monotonic behavior is lost. In this case, the initial evolution in transmissibility shortens the characteristic time scale of outbreaks, thus hampering immune escape. Moreover, under the transmission-recovery trade-off, viruses are able to find evolutionary stable strategies (ESS), thus eventually slowing down the evolution in transmissibility and therefore the chances for pathogens to become endemic. More future research will be relevant to determine how the phenomena here reported is sensitive to other biological trade-offs.
.4 Epidemiological and evolutionary time scales shape endemicity
Our results from stochastic simulations reveal that the interplay between immune pressure and evolution in transmissibility favors the endemicity of weakly infectious pathogens. To understand the origin of such phenomenon, we first realize that, in absence of transmissibility evolution (Fig. 2C), there is a critical value of the basic reproduction number ${\cal R}^{wt,C}_{0}(D_{x})$ below which extinction of the virus always occurs. This critical value is lower for viruses with efficient immune escape, i.e. high $D_{x}$ values. Inspecting visually the critical values in the endemicity curves for viruses now evolving their transmissibility (Fig. 3E), we notice that the unexpected non-monotonic behavior involves pathogens whose infectiousness falls below ${\cal R}^{wt,C}_{0}(D_{x})$. Conversely, the region ${\cal R}_{0}^{wt}>{\cal R}^{wt,C}_{0}(D_{x})$ retrieves the monotonically increasing behavior with the infectiousness of the wild-type variant. From this qualitative analysis, it becomes clear that the non-monotonic behavior previously described arises from the evolutionary dynamics undergone for viruses with ${\cal R}_{0}^{wt}<{\cal R}^{wt,C}_{0}(D_{x})$ to reach that critical value before becoming extinct.
To further support our intuition, we seek to analytically derive the eco-evolutionary constraints determining the endemicity of rapidly evolving viruses. Unfortunately, the underlying stochasticity of both evolutionary and epidemiological processes makes the analytical treatment of our model quite cumbersome, yet some theoretical results can be obtained in specific circumstances [32, 19]. To overcome this issue, we move to a more simplified eco-evolutionary framework, where we assume that both transmissibility and antigenic position evolve deterministically at a constant pace. Hence, for each time step $t$, there is a single variant circulating in the population, characterized by $x(t)=\widetilde{D}_{x}t$ and $\lambda(t)=\widetilde{D}_{\lambda}t$, where $\widetilde{D}_{m}$ represent the deterministic speed of evolution of trait $m$. Note that this framework cannot reproduce some of our previous findings, such as the accelerated evolution for the most innocuous viruses at initial stages of epidemic outbreaks. Notwithstanding its relevance, we will prove that such acceleration plays a minor role in the emergence of the non-monotonic behavior found for viral endemicity.
Introducing deterministic evolutionary pathways considerably eases the mathematical characterization of the underlying eco-evolutionary dynamics. In absence of evolution in transmissibility, we can assume that endemicity is determined by the prevalence of the disease at equilibrium, denoted in what follows by $\rho_{I}^{*}$. Using a travelling wave approach and considering a scenario with strong immune pressure, we show in the Materials and Methods section that the epidemic prevalence at equilibrium reads:
$\rho_{I}^{*}\simeq\frac{2\mathcal{R}^{wt}_{0}D_{x}}{\mu\pi}\ .$ (2)
If the prevalence $\rho_{I}^{*}$ is small enough, stochastic fluctuations associated with contagions and recoveries drive the system to the absorbing state and the virus becomes extinct. Considering the existence of a critical value $\rho_{I}^{*,C}$ to avoid the former scenario, we obtain that the minimum speed of evolution in the antigenic space that a virus requires to become endemic, denoted by $\widetilde{D}_{x}^{C}$, fulfills:
$\widetilde{D}_{x}^{C}=\frac{\mu a(N)}{\mathcal{R}^{wt}_{0}}\ ,$ (3)
where $a(N)=\pi\rho_{I}^{*,C}/2$ captures the aforementioned stochastic effects and depends on the population size $N$. In what follows, we will consider $a$ as a scaling factor to be fitted. To show the validity of this approach, we represent in Fig. 4A the fraction of realizations becoming endemic as a function of both $\widetilde{D}_{x}$ and $\mathcal{R}^{wt}_{0}$, proving that Eq. 3 captures the eco-evolutionary constraints allowing the persistence of viruses. As in the stochastic case, in absence of evolution in transmissibility, endemicity monotonically increases with $\mathcal{R}^{wt}_{0}$. Eq. 3 also captures the critical surfaces for pathogens with different $\mu$ values (SI Appendix, Fig. S3), showing that viruses with shorter infectious windows should evolve more quickly in their antigenic space to achieve endemicity. Likewise, we find that both $\widetilde{D}_{x}^{C}$ and, henceforth, $a$ are decreasing functions of the population size $N$, which is an expected result as stochastic fluctuations are less relevant in large populations (SI Appendix, Fig. S4).
Figure 4: Endemicity of viruses under deterministic evolution. A)-D): Fraction of endemic realizations (color code) as a function of the basic reproduction number of the wild-type variant $\mathcal{R}^{wt}_{0}$ and the speed of evolution in the antigenic space $\widetilde{D}_{x}$. The white solid line shows the theoretical estimation of the critical immunity escape value $\widetilde{D}^{C}_{x}$ delimiting the region $f_{endemic}=1$. Such quantity is obtained by setting $a=1/12$ and $b=5$ in Eq. 4. The values considered for the speed of evolution in transmissibility are: (A) $\widetilde{D}_{\lambda}=0$, (B) $\widetilde{D}_{\lambda}=4\cdot 10^{-5}$, (C) $\widetilde{D}_{\lambda}=8\cdot 10^{-5}$ and (D) $\widetilde{D}_{\lambda}=1.2\cdot 10^{-4}$. In all panels, the fraction of endemic realizations is obtained by performing 500 epidemic outbreaks and computing those persisting in the population after $t=1000$ days. The rest of model parameters are the same as in Fig. 2.
Now we modify Eq. 3 to capture how evolution in transmissibility shapes the endemicity of viruses. As the transmissibility evolves in an unbounded way, stochastic fluctuations during the epidemic bottleneck are the most relevant mechanism for the extinction of the virus. At that point, the virus infectiousness is determined by its gain in transmissibility during the first epidemic wave. Unfortunately, given the non-linearity of the SIR equations, the duration of such phase, denoted by $\tau_{wave}$, cannot be obtained analytically. Nonetheless, heuristically, we can argue that the characteristic time scale of an epidemic wave is proportional to the time to reach the peak of infected individuals, $t_{peak}$, which has been recently estimated very accurately [39]. Introducing the latter assumption into Eq. 3 yields:
$\widetilde{D}_{x}^{C}=\frac{\mu a(N)}{\mathcal{R}^{wt}_{0}+bD_{\lambda}t_{peak}(\mathcal{R}^{wt}_{0},\mu)k/\mu}\ ,$ (4)
where the expression for $t_{peak}$ is provided in the Materials and Methods section and $b$ is another scaling factor to be fitted to obtain the characteristic time scale of the disease, i.e. $\tau_{wave}=bt_{peak}$.
Figs. 4B-D show how the latter equation fairly captures the non-monotonic behavior of the endemicity with ${\mathcal{R}^{wt}_{0}}$ for different values of the speed of evolution in transmissibility $D_{\lambda}$. Note that there are some deviations for $\mathcal{R}_{0}\simeq 1$ and high $D_{\lambda}$ values, as the expression for $t_{peak}$ neglects the underlying evolution of the virus until reaching the epidemic peak, which is especially relevant in that region of the parameters space.
More importantly, Eq. 4 explains why viruses with intermediate infectiousness are less likely to become endemic. Namely, these viruses are not infectious enough to persist in the population without evolving their transmissibility but their generated first epidemic wave is not long enough to reach the critical transmissibility value ${\mathcal{R}_{0}^{wt,C}(D_{x})}$. Conversely, the nonlinear inverse dependence of $t_{peak}$ on $\mathcal{R}^{wt}_{0}$ (Eq. 11) makes the duration of the first epidemic wave much longer for pathogens with lower initial infectiousness, then giving them the chance to evolve before becoming extinct.
Discussion
Understanding the evolutionary forces governing how infectious diseases transition to become endemic in real populations represents a timely research question [40]. Here we propose a minimal eco-evolutionary framework to characterize how the evolution in both antigenic and non-antigenic traits shape the course and fate of epidemic outbreaks. Our results from stochastic simulations have first shown that evolution is not constant in both dimensions across time. Instead, the interplay between ecological and epidemiological processes accelerates the evolution in transmissibility at early stages of the outbreak while prioritizing the antigenic escape once the first epidemic wave is over. The accelerated evolution in fitness at early generations of RNA viruses has already been reported empirically in culture cells [41] and characterized theoretically through Fokker-Planck equations [42]. Likewise, the shift in the role of different evolutionary pathways has also been observed in the SARS-CoV-2 virus. Analysis of genomic data [43] reveals that intrinsic selection has become less relevant, with antigenic fitness gaining prominence throughout the COVID-19 pandemic.
Integrating the evolution of both antigenic and non-antigenic traits drastically changes the eco-evolutionary constraints shaping the endemicity of rapidly evolving viruses. Namely, we have observed how immune escape alone predicts that more infectious pathogens are more likely to become endemic. Conversely, in presence of evolution in transmissibility, viruses with low basic reproduction number, i.e ${\mathcal{R}^{wt}_{0}}\simeq 1$, are more likely to persist in the population that those with intermediate infectiousness, e.g. ${\mathcal{R}^{wt}_{0}}\simeq 2$. To characterize analytically this behavior, we have developed a simplistic model assuming that virus traits evolve deterministically. Our analysis reveals that the non-monotonic behavior emerges because the latter viruses generate much shorter epidemic outbreaks, thus not being able to evolve their transmissibility to reach the critical value needed to become endemic. Interestingly, the non-monotonic behavior of the endemicity of epidemic outbreaks has been already reported for other variants of the SIR model including demographic births and deaths [44, 45].
From a practical point of view, our findings represent a cautionary tale for the implementation of control strategies to manage epidemic scenarios. While the short-term beneficial effects of non-pharmaceutical interventions [46, 47] cannot be contested, determining their long-term impact on the modified epidemic trajectories still represents a more complex and intricate problem with growing attention in the scientific community [48, 49, 50]. Along this line, our results show that controlling an epidemic scenario by reducing the local effective reproduction number ${\cal R}$ might render unforeseen long-term consequences for disease control. In particular, our results show that mitigating an outbreak by flattening the epidemic curves induces a stronger selection pressure on the virus and increases the risk for their persistence in the population, due to the extended duration of the epidemic outbreak. Likewise, other works have reported that vaccination of the population also alters the antigenic space harboring the different variants of the virus [51, 52], thus potentially shaping the evolutionary strategies observed at the population level [53]. Moving from local to global interventions, the evolutionary fingerprint of travel bans and the subsequent geographical genetic isolation of different areas [54] remains to be solved; phylogeographic analyses [55] combined with multiscale epidemic frameworks [56] will be needed to address this question in future works.
Our findings should be considered in light of several limitations of the minimal framework here introduced. Biologically, the gradual evolution of the virus is challenged by the reported anomalous diffusion behaviors of viruses across the genomic space [57] or the complex genotype-phenotype networks shaping the evolution of viruses [58]. The evolutionary dynamics in our model could also be refined by introducing trade-off mechanisms [59], coupling the evolution of different epidemiological parameters. Despite the scarce empirical evidence supporting them [60], evolutionary trade-offs are typically included in models to warrant the existence of evolutionary stable strategies [61]. Our first exploration with the recovery-transmissibility trade-off shows an uneven impact of such mechanism, keeping our findings for some viruses but altering them for others. At the population level, assuming all the individuals to be equivalent does not allow our model to capture the strong influence reported for superspreaders [62, 63] or immuno-compromised patients [33, 64] on the eco-evolutionary dynamics of viruses. Notwithstanding all these limitations, the generality of our findings, based on the relationship between epidemiological and evolutionary time scales, make us feel confident about their ubiquity in more biologically-grounded models.
In a nutshell, our results underscore the relevance of the interplay between different evolutionary pathways of rapidly evolving virus in shaping their associated epidemic trajectories. Improving our knowledge on this topic will enhance our preparedness against future epidemic threats and the use of eco-evolutionary frameworks as reliable benchmarks to predict the evolution of emergent pathogens and design optimal control policies to mitigate both their short-term and long-term impact on society.
Materials and Methods
Stochastic simulations
To obtain the epidemic trajectories shown in the manuscript, we perform synchronous discrete-time stochastic simulations using the extended version of the SIR model described in Fig. 1, assuming $\Delta t=1$. We split each time step into two different stages governing epidemiological processes and virus evolution respectively.
At the first stage, we assume that infected individuals recover with a probability $Pr(I\rightarrow R)=1-e^{-\mu\Delta t}$. To simulate contagion processes, we assume that each susceptible agent $i$ chooses $k$ neighbors at random and compute the probability of infection as $Pr(S_{i}\rightarrow I)=1-e^{-\sum\limits_{l\in\Gamma(i)}\lambda_{l}\Delta t}$, where $\Gamma(i)$ denotes the set of infected neighbors of agent $i$ at that time step. If a contagion event is successful, the variant contracted by the agent is chosen proportionally to the infection rates of the variants existing in $\Gamma(i)$. Therefore, the probability that agent $i$ contracts the disease from an infected neighboring agent $j$ is $Pr(S_{i}\rightarrow I_{j})=\dfrac{\lambda_{j}}{\sum\limits_{l\in\Gamma(i)}\lambda_{l}}Pr(S_{i}\rightarrow I)$. For reinfection processes, the same rationale to choose the variant applies but now considering that the individual rates at which a recovered agent $i$ contracts the disease from an infected agent $j$ is $\lambda^{\prime}_{ij}=\Theta(x_{j}-x_{i})\lambda_{j}\left(1-e^{-(x_{j}-x_{i})}\right)$, where $\Theta(x)$ is the Heavyside function introduced to avoid reinfection from past variants.
Once the epidemiological state of each individual is updated, evolution of the virus only occurs inside infected individuals. Therefore, the transmissibility (antigenic position) of an agent $j$ variant $\lambda_{j}$ ($x_{j}$) changes according to $\lambda_{j}(t+\Delta t)=\lambda_{j}+D_{\lambda}\Delta t\mathcal{N}(0,1)$ $\left(x_{j}(t+\Delta t)=x_{j}+D_{x}\Delta t\mathcal{N}(0,1)\right)$, where $\mathcal{N}(0,1)$ represents a random number drawn from the standard normal distribution.
Approximation to the case reproduction number $\mathcal{R}_{case}(t)$
In the context of an epidemic outbreak, $\mathcal{R}_{case}(t)$ computes the expected number of contagions made by individuals becoming infected at time $t$. The exact computation of such quantity requires knowing the shape of the distribution of variants in both the transmissibility and antigenic spaces. For the sake of simplicity, we opt for providing a rough estimation of the case reproduction number, $\mathcal{R}^{app}_{case}(t)$. For such purpose, we assume that the standard deviation of the aforementioned distribution $\sigma_{x}$ represents the typical distance observed in the antigenic space and that the average transmissibility $\langle\lambda\rangle$ also constitutes a representative quantity for contagion dynamics. Moreover, let us also neglect the changes in the pool of susceptible and recovered population and the evolution of the virus during the infectious period of the focal agent. Under these assumptions, the approximated case reproduction number can be readily computed as:
$\mathcal{R}^{app}_{case}(t)=\frac{\langle\lambda\rangle k}{\mu}\left[\rho_{S}(t)+\rho_{R}(t)\left(1-e^{-\sigma_{x}}\right)\right]\ ,$ (5)
where $\rho_{m}(t)$ represents the fraction of population in state $m$ at time $t$.
Epidemic prevalence at equilibrium for a virus with deterministic antigenic escape
In this section, we derive the epidemic prevalence at the endemic equilibrium for a virus evolving deterministically its antigenic position and not evolving its transmissibility. For such purpose, let us first define $\rho_{I}(x,t)$ ($\rho_{R}(x,t)$) as the fraction of the population in the infected (recovered) state carrying the variant $x$ at time $t$. Considering antigenic escape and reinfection events, the time evolution of the latter is given by [19, 32]:
$\frac{d\rho_{R}(x,t)}{dt}$ $=\mu\rho_{I}(x,t)$
$-\lambda k\rho_{R}(x,t)\int\rho_{I}(x^{\prime},t)\left(1-e^{-(x^{\prime}-x)}dx^{\prime}\right)$ (6)
Assuming a deterministic evolution of the virus greatly simplifies the set of equations governing the unfolding of the epidemic trajectories, as there is a single circulating strain in the population. Mathematically, this implies that $x_{c}(t)=\widetilde{D}_{x}t$ for the circulating pathogen and hence $\rho_{I}(x,t)=\rho_{I}(t)\delta\left(x-x_{c}(t)\right)$, where $\delta$ represents the Dirac delta function. Rewriting Eq. 7 using a travelling wave approach, we obtain:
$\frac{d\rho_{R}(\tau)}{d\tau}=-\lambda k\rho_{R}(\tau)\rho_{I}(0)\left(1-e^{-\widetilde{D}_{x}\tau}\right)\ ,$ (7)
where $\tau$ refers to the time elapsed since the recovery of the agent. Assuming that the disease has reached the endemic equilibrium, i.e. $\rho_{I}(0)=\rho_{I}^{*}$, and a regime of strong immune pressure, i.e. $\widetilde{D}_{x}\tau\ll 1$, we solve the previous ODE obtaining:
$\rho_{R}(\tau)=\mu\rho_{I}^{*}e^{-\widetilde{D}_{x}\lambda k\rho_{I}^{*}\tau^{2}/2}\ ,$ (8)
[…]
$\sqrt{\frac{\mu\pi\rho_{I}^{*}}{2\mathcal{R}_{0}\widetilde{D}_{x}}}+\rho_{I}^{*}=1\ .$ (9)
Assuming a very small prevalence in the endemic regime, i.e. $\rho_{I}^{*}\ll 1$, we can estimate $\rho_{I}^{*}$ as
$\rho_{I}^{*}\simeq\frac{2\mathcal{R}_{0}\widetilde{D}_{x}}{\mu\pi}\ .$ (10)
Time for an epidemic peak in a SIR model
For a well-mixed population with constant size, the exact time to the epidemic peak $t_{peak}$ in the standard SIR model can be approximated as [39]:
$t_{peak}(\mathcal{R}^{wt}_{0},\mu,i_{0})=\frac{1}{\mu}\frac{1}{\mathcal{R}^{wt}_{0}-1}\ln\left(\mathcal{R}^{wt}_{0}+i_{0}^{-1}\left(\mathcal{R}^{wt}_{0}-1\right)^{2}\right)$ (11)
where $i_{0}$ refers to the fraction of population initially infected. This expression represents a first-order approximation to the exact computation of the epidemic peak in the standard SIR model. Nonetheless, it provides reasonably accurate estimations of the peak in most of the epidemic scenarios [39]. In our model, the latter expression is used to obtain a characteristic time scale during which viruses can evolve their transmissibility to avoid getting extinct. Consequently, note that the former expression comes with intrinsic limitations, as it assumes a fixed value of the infectiousness $\lambda$ and not a evolving quantity over time. Therefore, we should expect predictions to be misaligned with the actual epidemic peak for quickly evolving viruses with low initial infectiousness, i.e. ${\mathcal{R}_{0}\simeq 1}$. In that region, the slow propagation of the virus leads to a substantial evolution in the virus infectiousness throughout the first epidemic wave, which in turn produces a much shorter epidemic wave than the one predicted by Eq. 11.
Acknowledgements.
The author thanks A. Alsina, S. Lamata-Otín, J. Gómez-Gardeñes and A. Arenas for their extremely valuable feedback on the results here discussed and their careful reading of the first version of the manuscript. The author also thanks L. Arola, R. Vishwakarma, P. Castioni and L. Fant for insightful discussions about the paper. The author acknowledges financial support through grants JDC2022-048339-I and PID2021-128005NB- C21 funded by MCIN/AEI/10.13039/501100011033 and the European Union “NextGenerationEU”/PRTR”.
Supplementary Information: Eco-evolutionary constraints for the endemicity of rapidly evolving viruses
Stochastic evolution of antigenic and non-antigenic traits
In this section, we extend the results reported in Figure 3 about the influence of the stochastic evolution of both antigenic and non-antigenic traits on epidemic outbreaks. We start by considering a fixed value of the speed of evolution in antigenic escape ($D_{x}=0.0003$) and different speeds of evolution in transmissibility $D_{\lambda}$. First, we can observe in Fig. S1A how increasing the speed of evolution shrinks the duration of the first epidemic wave and reduces the convergence time to the endemic phase. Regarding the virus traits, Fig S1B confirms that the evolution in transmissibility is accelerated at early stages of the outbreak for pathogens with low infectiousness. As stated in the main manuscript, this acceleration is more pronounced for high speeds of evolution, giving rise to an earlier collapse of the curves associated to different infectiousness of the wild-type variants. Analogously, Figs. S1C illustrates that, regardless of the speed of evolution in transmissibility, most antigenic variation is produced at later stages of the outbreak. Again, we find how quickly evolving viruses tend to collapse to the same trajectories in the antigenic space earlier. Finally, we are interested in determining how the evolution in transmissibility shapes the dependence of virus endemicity on the reproduction number of the wild-type variant. Fig. S1 shows that, for $D_{x}=0.0003$, the non-monotonic behavior penalizing the endemicity of viruses with intermediate infectiousness appears consistently regardless of the speed of evolution in transmissibility. When the immune pressure is weaker, we instead observe how the non-monotonic behavior disappears as viruses evolve more quickly, yielding the usual increase in endemicity with the infectiousness of the wild-type variant. These results show that changing their intrinsic transmissibility represents an alternative evolutionary pathway for viruses to partially overcome the immune pressure existing in the population.
Transmission-recovery trade-off
The model presented in the main manuscript does not include any biological trade-off linking the different epidemiological parameters of pathogens and, therefore, their evolution. These trade-offs are widely adopted in the modelling community to produce evolutionary stable strategies (ESS) bounding the evolution of viruses. Among them, the virulence-transmissibility trade-off is undoubtely the most studied one following the seminal work by Anderson & May [1]. The virulence-transmissibility trade-off assumes that the virulence of a virus, defined as the rate at which the host die because of the pathogen, increases with the virus transmissibility. Such increased lethality shortens the infectious period during which the virus host can transmit the pathogen to other individuals. This double-edge sword creates an optimal strategy at intermediate infectiousness, maximizing the expected size of the offspring produced by infected individuals in next generations.
As our framework does not incorporate any birth-death processes involving hosts, we opt for not including the virulence-transmissibility trade-off. Instead, we can study how the joint evolution of the transmissibility and the infectious period affect the eco-evolutionary dynamics of pathogens by introducing the so-called transmission-recovery trade-off [2]. This trade-off implies that the infectious period is shortened for highly transmissible pathogens and emerges naturally when both the production of immune defenses against a pathogen and its inter-host transmissibility are assumed to be proportional to the pathogen load inside hosts.
To account for the transmission-recovery trade-off, we decompose the recovery rate of individuals $\mu$ into two components: the baseline recovery rate $\mu_{0}$, constant and independent of the pathogen load, and the recovery rate $\mu^{\prime}$ changing over time as a result of the evolution in transmissibility $\lambda$ and the aforementioned trade-off. Mathematically, we assume $\lambda=\gamma\sqrt{\mu^{\prime}}$, where $\gamma$ is a constant factor related to the basic reproduction number of the wild-type variant ${\cal R}^{wt}_{0}$ and the initial value of the evolving recovery rate $\mu^{\prime}_{(t=0)}$. Namely:
$\gamma=\frac{{\cal R}_{0}^{wt}\left(\mu_{0}+\mu^{\prime}_{(t=0)}\right)}{k\sqrt{\mu^{\prime}_{(t=0)}}}\ .$ (S1)
To study the relevance of the transmission-recovery trade-off, we fix the speed of evolution in tranmissibility to $D_{\lambda}=0.0003$ and the immune escape to $D_{x}=0.015$ and study how epidemic trajectories depend on the initial value of the evolving recovery rate $\mu^{\prime}$. When the transmission-recovery trade-off plays a minor role, i.e when $\mu^{\prime}_{(t=0)}\ll\mu_{0}$, the epidemic curves resemble those shown in Fig.3, characterized by a first prominent epidemic wave followed by an epidemic bottleneck and an endemic regime with a steady growth rate in the number of cases. Nonetheless, as both rates get comparable, the trade-off bounds the evolution of the virus and slows down considerably the growth rate of cases in the endemic regime, yielding a roughly constant prevalence of the disease for some epidemic scenarios. This deceleration is better illustrated in Fig. S2B, as the basic reproduction number ${\mathcal{R}_{0}}$ converges to the corresponding ESS as the trade-off gains more relevance.
Regarding the impact on immune escape, the trade-off forces the variants to be more widespread across the antigenic space, as shown in Fig. S2C. This phenomenon is explained by recalling that evolution increases transmissibility but also shrinks the infectious period of agents, which prevents viruses from overcoming the immunity within the population unless more antigenic variation is accumulated. Finally, we show in Fig. S2D that the non-monotonic behavior of the endemicity of viruses on the reproduction number of the wild-type variant is hindered as the transmission-recovery trade-off becomes more relevant. We believe this phenomenon arises from the shortening of the infectious period and the slowing down in the evolution of transmissibility, making it difficult for weakly infectious pathogens to reach the critical values of transmissibility needed to become endemic.
Endemicity of viruses with deterministic evolution
In the Materials and Methods section of the main manuscript, we derive a theoretical expression for the minimum immune escape rate $\widetilde{D}^{C}_{x}$ that warrants the stability of the endemic equilibrium, yielding:
$\widetilde{D}^{C}_{x}=\frac{\mu a(N)}{\mathcal{R}^{wt}_{0}}\ ,$ (S2)
where $a(N)$ is a function of the population size and encompasses the stochastic effects destabilizing the endemic equilibrium and moving it to the absorbing state.
While in the main manuscript we focus on showing the inverse dependence of $\widetilde{D}^{C}_{x}$ on $\mathcal{R}^{wt}_{0}$, here we are interested in exploring how the boundary of the endemic regime depends on the rest of parameters. To address this question, we first set the population size to $N=10000$, consequently fixing the constant $a=1/6$ (Fig. 4A), and represent in Fig. S3 the fraction of endemic realizations as a function of the reproduction number ${\cal R}_{0}^{wt}$ and the immune escape $\widetilde{D}_{x}$ for three values of the duration of the infectious window $\mu^{-1}$: $\mu^{-1}=10$ days (Fig. S3A), $\mu^{-1}=9$ days (Fig. S3B), $\mu^{-1}=8$ days (Fig. S3C). There we check that the theoretical estimation provided by Eq. S2 fairly captures the endemic regime for the three epidemic scenarios and confirm that the critical value $\widetilde{D}^{C}_{x}$ increases as the infectious period is reduced.
To estimate finite size effects on the virus endemicity, we represent in Figure S4 how the fraction of endemic realizations changes a function of the immune escape $\widetilde{D}_{x}$ and the population size $N$ for two pathogens differing in their reproduction number: $\mathcal{R}_{0}^{wt}=3$ (Fig. S4A) and $\mathcal{R}_{0}^{wt}=4$ (Fig. S4B). In both panels, we clearly show how the critical immune escape for endemicity decreases as populations get larger, thus encoding an inverse relationship between the factor $a$ and the population size $N$. Interestingly, the curves seems to collapse for large population sizes ($N\simeq 20000$), thus allowing the qualitative findings reported in the manuscript to be extrapolated to other greater population sizes.
[…]
Figure S3: Endemicity of viruses under deterministic evolution. A)-C): Fraction of endemic realizations (color code) as a function of the basic reproduction number of the wild-type variant $\mathcal{R}^{wt}_{0}$ and the speed of evolution in the antigenic space $\widetilde{D}_{x}$. The white solid line shows the theoretical estimation of the critical immunity escape value $\widetilde{D}^{C}_{x}$ delimiting the region $f_{endemic}=1$. Such quantity is obtained by setting $a=1/12$ and $b=5$ in Eq. 4 of the main text. The values considered for the duration of the infectious windows are: (A) $\mu^{-1}=10$ days, (B) $\mu^{-1}=9$ days, (C) $\mu^{-1}=8$ days. In all panels, the speed of evolution in transmissibility is set to $\widetilde{D}_{\lambda}=0$. Moreover, the fraction of endemic realizations is computed by performing 500 epidemic outbreaks and computing those persisting in the population after $t=1000$ days. The rest of model parameters are the same as in Fig. 2 of the main manuscript.
Figure S4: Finite size effects on the endemicity of viruses under deterministic evolution. Endemicity $f_{endemic}$ of the virus as a function of the basic reproduction number of the speed of evolution in the antigenic space $\widetilde{D}_{x}$ and the population size $N$ (color code). Two different values for the reproduction number of the wild-type variant are considered: (A) $\mathcal{R}_{0}^{wt}=3$ and (B) $\mathcal{R}_{0}^{wt}=4$. In all panels, the speed of evolution in transmissibility is set to $\widetilde{D}_{\lambda}=0$ and the rest of epidemiological parameters are the same as in Fig. 2 of the main manuscript.